Marks: 2
Solve the system of equations by the best way you think is possible. Put the answer in an ordered pair.
2x - 6y = 6
2x + 3y = 24

Found 2 solutions by sofiyac, ewatrrr:
Answer by sofiyac(983)

2x - 6y = 6
2x + 3y = 24 multiply second one by 2 and add to first
**************
2x - 6y = 6
4x + 6y = 48
*************
6x=54 divide each side by 6
x=9 plug this into either of the original equations and solve for y
2*9-6y=6
18-6y=6
-6y=-12
y=2
so the solution is (9,2)

Answer by ewatrrr(24785)

 
Hi
Solving using the elimination Method
2x - 6y = 6 Green
2x + 3y = 24 Blue
-2x + 6y = -6 ||Multiplying 1st EQ thru by -1 to eliminate x
2x + 3y = 24
9y = 18
y = 2 and x = 9 ( 2x - 6y = 6, x = 18/2)
Ordered pair (9,2) is the solution for this system